Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used iRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the predetermined overhead rate for the year. Predetermined ovRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the amount of underapplied or overapplied overhead for the yeaRequired 1 Required 2 Required 3 Required 4 Required 5 Prepare a schedule of cost of goods manufactured for the year. AssumeCompute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answeAssume that the $35,000 ending balance in Work in Process includes $8,100 of direct materials. Given this assumption, supply
Answer #1

1) Predetermine OH rate = 139500/93000 = 150% of direct material used

2) Direct material used = 28000+132000-13000 = 147000

Applied overhead = 147000*1.5 = 220500

Actual overhead = 146100+8500+20000+11000+7300+31000 = 223900

Under applied overhead = 223900-220500 = 3400

3) Schedule of cost of goods manufactured

Direct materials
Beginning raw material 28000
Raw material purchase 132000
Raw material available 160000
Less: Ending raw material -13000
Raw material used in production 147000
Direct labor 89000
Applied overhead 220500
Total manufacturing costs 456500
Beginning work in process 50000
Total Cost of work in process 506500
Less: Ending work in process -35000
Cost of goods manufactured 471500

4) Unadjusted cost of goods sold = 68000+471500-61000 = 478500

5) Calculate following

Direct material 8100
Direct labor 14750
Manufacturing overhead (8100*1.5) 12150
Work in process inventory 35000